Gundam 0083 Stardust Memory - Episode 11: La Vienne Rose Page 1

Gundam 0083 Stardust Memory Episode 11: La Vienne Rose

Translated By Tonghyun “Vajra” Kim (tkim@sdcc3.ucsd.edu) Edited by Noel Gamboa (egamboa@sdcc13.ucsd.edu) OPERATION X Version 0.9 9/22/92

Narration

The bay of Solomon was filled with tragedy and blinding lights. Disgrace that burns the eyes… Borrowing the power of Unit 1, Kou took on Unit 2. They were drawn to each other as if they were trying to put an end to their destiny. The battle went on without end. But Uraki failed to realize that this was naught but the prologue to Operation Stardust. When both Gundams perished to the Sea of Solomon, Operation Stardust finally revealed its true nature.

(Cima’s fleet, escorting the hijacked colony, joins up with the Delaz Fleet.)

Helmsman

His Excellency Delaz’s fleet is approaching from portside. We will be joining them.

Delaz fleet bridge officer

All ships, begin resupplying the Cima Fleet.

Delaz

I can picture the panic going on in Jaburo. What kind of reply will the Feds come up with? No matter how they respond, the awakening Stardust cannot be stopped. The arrow has already been fired.

(Federation officials and staff generals meet in Jaburo to discuss their course of action.)

Staff 1

Nonsense! Does Delaz seriously think that we’ll accept this kind of demand?

Staff 2

Even if they drop a colony on the moon, The Federation won’t budge an inch.

Staff 3

Whatever’s on his mind, we should make sure that it doesn’t have a bad effect on the political situation.

Staff 4

So you want to stay in power?

Staff 5

We just want to emphasize that we all share the same view on the importance of the moon.

Staff 6

So this is the so-called Civilian Control.

Staff 7

By the way, what should we do with a new meddler in the game, Axis?

Staff 8

If they remain neutral, we’ll do ’em a favor and just leave them alone. There’s no need to make them our enemy.

Staff 9

Look. Here is the colony that’s heading for the moon. And this is the remaining fleet from Konpei. At full speed, the fleet can intercept the enemy just before the colony hits the moon, then destroy the enemy force, and ignite the thruster fuel, therefore foiling the enemy’s plan.

Staff 10

And we’ll be done exterminating pirates.

Staff 5

But I still don’t understand. Why didn’t they use the thruster fuel when they jacked the colony, rather than making two colonies collide with each other?

Staff 11

Simply put, they didn’t have the energy to ignite it.

Staff 12

I love the calm sky.

[AXIS ADVANCE FLEET]

Earlier, the Axis Fleet joined with the Gato Fleet [a branch of the Delaz Fleet headed by Gradol’s Sanzibal-class ship]. Gato is onboard the Axis flagship, speaking with its captain.

Hustler

I’m truly impressed by your fight. we really want to help you but…

Gato

We’re always thankful for your support. Thanks to Axis Fleet, who are neutral, we can fight without worrying about the aftermath. Please retrieve the soldiers after everything comes to an end.

Hustler

I give you my word.

Hustler

Take this as our departing gift. This is the least we can do.

Gato

That’s… Magnificent!

Gato

As if the spirit of Jion itself has taken physical form.

Hustler

This is the mobile armor Neue Ziel. (Neue Ziel: German for new goal)

The Albion stops at La Vienne Rose.

Scott

Approaching La Vienne Rose. Distance 8000.

Basrof

Activating laser sensor.

Operator

This is La Vienne Rose. We’ll begin guiding you.

Maurice

Laser catch. Calibrating the pulse now. Ten minutes to locking.

Operator

Guide beacon normal. Relative velocity: 11.

Inside the Albion, en route to the main bridge…

Keith

Hey, Kou!

Kou

Keith.

Keith

At last, we’re gonna get Unit 3.

Kou

Yeah…

Keith

What’s wrong? You look depressed. We’re going ashore!

Kou

Nina… is still against my riding Unit 3.

Sinapus

Begin the mooring process.

Basrof

Roger.

Sinapus

Nina. As I expected from the largest ship of Anaheim. It seems perfectly suited for testing Unit 3.

Nina

Yes…

Scott

Be ready to turn up immediately.

Keith

Kou, what’s Unit 3 like, anyway?

Kou

Only the Captain and Nina know the details. It’s said to be the unit for the defense of the strategic point in space.

Keith

For the defense of the strategic point in space? What’s that?

Kou

I’ll make sure I can pilot it, no matter what kind of unit it is.

Kou and Keith walk into the bridge.

Sinapus

Lt. Uraki.

Kou

Lt. Uraki, reporting for duty.

Nakato

He’s the Gundam pilot?

Sinapus

That’s right.

Nakato puts on a cynical smile.

Nakato

You are?

Nakato

I’m Commander Nakahha Nakato of Security Squadron 401. I’ll inform you of the situation since you are the Gundam pilot. The Albion will be under our squadron’s supervision during its stay on La Vienne Rose. The Albion has been relieved from its mission of pursuing the Delaz Fleet. Your new duty is to guard the La Vienne Rose and Gundam Unit 3.

Keith

Hold on a second, sir. We need to get Unit 3 quickly and stop the falling colony…

Nakato

The pursuit fleet from Konpei will take care of the colony.

Keith

But we’re on this mission with General Kowen’s approval…

Nakato

The Gundam Development Plan has been removed from Gen. Kowen’s hands. Do I have to tell you why? You lost Unit 2 and let the Naval Review get totally trashed! And…

Sinapus

Commander, I’d like to take Lt. Uraki onboard for the transfer of combat data.

Nakato

That’s fine with me. Only the Captain, Nina Purpleton, and he will be allowed to go ashore.

Delaz

Can’t the Feds do any better than give out such a naive reply? We’ve already accomplished our goal, with the Federation approving the presence of Axis Fleet.

Delaz

Is Cima carrying out her part?

Delaz fleet bridge officer

Aye, sir. Everything’s proceeding as scheduled.

Clehna

Welcome to La Vienne Rose. I’m Clehna Hacswell, the chief here.

Sinapus

I’m Sinapus, captain of the Albion.

Kou

I’m Lieutenant Junior Grade Uraki.

Nina

It’s been so long since I’ve seen you, Chief.

Clehna

Oh dear, Nina.

Clehna

I wish that the others could come onboard La Vienne Rose. I’m sorry about that.

Clehna

Over here please. I believe you’ve already met Commander Nakato.

Nakato

Sorry for the inconvenience.

Clehna

And she’s the one who’s in charge of Unit 3.

Lucette

It’s a pleasure to meet you. I’m Lucette Audevie. Long time no see, Nina.

Nina

You look well, Lucette.

Lucette

I heard about Unit 1. How terrible!

Nina

Yes…

Lucette

I heard a lot about you, Lt. Uraki. I heard that you’re also well versed in the mobile suit theories. Please test pilot Unit 3…

Nakato

Ms. Audevie. The project’s been suspended. And I can’t let him pilot Unit 3.

Lucette

Why? Unit 3’s already completed. The only thing left is to collect the data for real combat!

Nakato

I’m a soldier. I have to obey orders.

Lucette and a very quiet Nina, at La Vienne Rose’s mess hall…

Lucette

It’s actually the end of the Development Plan! I have no idea what the military has in mind. Commander Nakato replaced the test staff as soon as he arrived here.

Lucette

I think that you, who developed Unit 1… can understand how I feel, right?

Nina

Yes, I do.

Nina

But the research can’t go on without the test pilot, can it?

Lucette

I can’t accept this! Unit 3 is perfect! When it’s completed… Say, can I borrow Lt. Uraki?

Nakato

Kou?

Lucette

His ability to make split-second decisions, his intuition… and above all, he pilots with an understanding of the mechanism.

Lucette

I believe he can handle even the complex weapon control system of Unit 3.

Nakato

No way! Never!

Lucette

Why are you so upset about that?

Nina

It’s already too late for the battle anyway…

Lucette

Nina, something must’ve snapped in you after you lost Unit 1.

A squadron of Federation ships spots the hijacked colony.

Operator

I can see it! The colony!

Ship captain

Inform the main fleet of this. We’ve spotted the enemy fleet.

Operator

Aye, sir!

Ship captain

We’re gonna do it. All hands, battle stations - bombardment!

Operator

A high-level heat source has been detected. It’s huge!

Ship captain

From above? What is it?!

The Neue Ziel rockets towards the five Federation vessels.

Gato

Meet your doom!

The four ships experience sudden death. Meanwhile, Kou walks out of a shower seconds before the doors to his quarters opens.

Kou

Damn! That Commander Nakato…

Lucette

I’m sorry.

Kou

Lucette… No, it’s okay.

Lucette

Uh… can I bother you for a moment?

Lucette

I’ve something to show to you, Lieutenant.

Kou

To me? Is it about Unit 3?

Lucette

Take a look at this.

Lucette pops up some Unit 3 stats and graphics on a laptop.

Kou

Unit 3, I knew it!

Lucette

Unfortunately, most of the data is in the hands of Commander Nakato and this is all I have.

Kou

I can’t believe this. These are all…

Lucette

Well? Think you can handle it? It’s up to you, whether you get Unit 3 or not.

Kou

But that’s against the order…

Lucette

Don’t you need Unit 3 right now? Why did you come here? Even the brass will understand since we’re doing this for the Federation.

Lucette

Please excuse me for now. I’ll be expecting a good answer from you.

As Lucette walks out, the doors open. Nina is standing outside.

Kou

Nina.

Lucette

Nina, the Lieutenant accepted my offer.

Nina

Is it true, Kou?

Kou

Nina. I…

Nina

Don’t! Why do you have to go that far?

Kou

We don’t have time. While we’re wasting time here, Gato and his people are…

Kou

I have to change clothes.

Kou shuts the door. On the moon, Anaheim’s vice-CEO and presumably some lunar government officials listen to Cima’s voice.

Cima

My sympathies to you. You’re running out of time. It’s already en route to collide with the moon. Even your last hope, the Fed’s fleet, is powerless before Gato.

Operator

Begin adjusting the mirror’s angle!

Operator

Activate number one. Activate number three!

Operator

Open the detonation condenser’s circuit!

Cima

We have no wish to drop a colony on the moon. That’s why I’m telling you how to save Von Braun. You have only two choices. To be smashed to death by the falling colony… or…

Beams shoot out from the moon, hitting the colony. Seconds later there is a flash of light and a thundering explosion.

Operator

The colony has gained thrusting force. It’s begun accelerating.

Base commander

What? The propellants for repositioning the colony have been ignited? Where is it headed now?

Operator

The colony which gained thrusting force… will desert the lunar orbit, and this time… it’ll be drawn by the earth’s gravity.

Base commander

We’ve been tricked!

Kou

A feint? The colony drop’s real target is the earth!?

Lucette

It’ll bring back the tragedy of the One Year War. There’s only 38 hours until it collides with the earth.

Kou

But why the feint… ? I see! To make the pursuing fleet waste their propellants!

Lucette

There’s no time to hesitate. Lt. Uraki!

One by one, ships from the Konpei pursuit fleet drop out.

Ship 1

This is Madros. We ran out of propellant! We’re slowing down.

Ship 2

This is Mongrel. We’re losing thruster force. We have to drop behind the battle line.

Ship 3

Damn! How many ships can stay if this keeps up?

Base commander

When will the supply ship arrive?

There’s the sound of a brief scuffle, followed by shadows entering a dark hangar.

Kou

Lucette!

Lucette

Lieutenant, over here!

Military police officer

Commander! The Albion pilot has entered the factory block!

Nakato

What? Drag him out! Don’t let him touch Unit 3! I authorize you to open fire if he resists!

Nakato looks at Sinapus, and sees the Captain staring icily back at him.

Nakato

Shit! Let’s go!

Nina looks pleadingly towards Captain Sinapus, who takes a communicator from his pocket. Inside the hangar, Kou has slipped into an Anaheim Electronics pilot uniform.

Kou

Lucette, I’m ready for it anytime.

Lucette

Nice timing, Mr. Lieutenant.

Hangar lights unveil Unit 3.

Kou

This isn’t the Core Block System… And that Tail Binder can also be used as a Locking System.

Lucette

You passed the test, Lieutenant. Hurry, I’ll fill you in on the details later.

Nakato

That’s far enough, Uraki!

Nakato

What the hell do you think you’re doing, Lieutenant? Move away from the Gundam!

Nakato draws his pistol and aims squarely at Uraki.

Nakato

This is not just a threat!

Lucette

Lt. Uraki, get in the cockpit quick!

Nakato

Don’t test my patience!

Nakato fires at Kou’s feet.

Nakato

If you surrender now, I’ll let the case go unnoticed.

Kou

Damn you… Can’t you see what this is all about? The colony - !

Kou, perhaps realizing that his words are wasted on Nakato, seals his helmet shut and turns around.

Nakato

You fool!

Nakato fires.

Lucette

Lt. Uraki!

Lucette jumps into the line of fire.

Kou

Lucette!

Nakato

You bastard, you used a civilian as your shield! You tried to steal Unit 3 and you murdered a civilian. You should get a court martial. But I’ll have the pleasure of executing you myself.

Machine-gun fire cuts through the air!

Bait

You pig-headed morons! Drop your guns right now!

Monsha

Hey Commander! Aren’t you playing a little too rough?

Nakato

You bastards! When did you…?

Sinapus levels a pistol at Nakato’s skull.

Sinapus

Withdraw your men, Commander Nakato. Our enemy is the Delaz Fleet.

Nina

Lucette!

Kou

Nina! Take care of her! A doctor! Bring a doctor quick!

Nina

Lucette.

Lucette

Nina… please take care of Unit 3. I… knew it all along… why you were so strongly opposed to Unit 3.

Nina

You knew about Kou and me?

Lucette

It’s not that, not just because you’re worried about his safety, but also because of what you’re most afraid of…

Nina

Lucette… so you…

Mora

Doctor, hurry!

Lucette

Three years ago… Because I knew that, three years ago on the moon… you and Lt. Gato… were lovers.

Nina

Because… Because I didn’t want their fight to go on anymore…

Lucette

Nina… I know that you can… take care of Unit 3… Unit 3…

Nina

Lucette… Lucette!

The Albion leaves La Vienne Rose.

Sinapus

May God have mercy on Lucette Audevie’s soul.

Nina and Mora ready Unit 3 for launch. Kou is getting a crash course in piloting Unit 3.

Mora

They call this a mobile suit!?

Kou

Damn! Even with auto mode on, my timing’s off…

Nina

Watch the evaluation monitor. You must finish loading the weapon container inside the armor in three seconds.

Gato continues to smash any and all Federation ships that attempt to follow the hijacked colony.

Gato

I guess that was the last of the enemies who followed us. A new enemy? But… there is no data on it.

Dendrobium Stamen, the combination of Gundam Unit 3 Stamen and its armed base Orchis, rockets toward Neue Ziel!

Preview

The Critical Defense Point, the last place where the falling colony can be intercepted. The Critical Defense Point, the place where everything will end. Beyond this place, people will not change destiny. Gundam Unit 3 and Neue Ziel, fateful Uraki and Gato… Not much time left until the moment when their fates are decided. Little time left until the colony strikes the earth.

Next: Critical Defense Point

END